determine whether the condition is true or false. assume a =2 and b =3

(5-a)*b<7

a^b =b^a

3E-02 < .01*a

(a*a< b) Or Not (a*a < a)

Not (a < b) Or Not (a < (b +a))

(( a=b) Or Not (b < a)) And ((a , b) Or (b = a +1))
Answer #1
  1. (5-2)*3<7 ==> 3*3<7 ==> 9<7 ==> False
  2. 2^3=3^2 ==> 8=9 ==>False
  3. 3E-02 < 0.01*2 ==> 3*10^-2 < 0.02 ==> 0.03<0.02 ==> False
  4. (2*2<3) Or !(2*2<2) ==> (4<3) Or !(4<2) ==> False Or not False ==> False Or True ==> True
  5. !(2<3) Or !(2<(2+3)) ==> not true Or not(2<5) ==> False Or not true ==> False Or False ==> False
  6. ((2=3) Or !(3<2)) And ((2,3) Or (3=(2+1)) ==> (False Or not False) And (( 2,3) Or (3=3)) ==> (False Or True) And ((2,3) Or True) ==> True And True ==> True